### v1

@@ 2.4 Correlation between DFT and magnetic Moment Tensor Potential equilibrium magnetic moments
For the additional validation of the fitted ensembles of potentials we estimate the quality of atomic magnetic moment equilibration. For this purpose we choose the configurations with the equilibrium magnetic moments from the training set, equilibrate these magnetic moments with the ensembles of mMTPs, and compare the result with DFT. The correlation between the DFT and mMTPs equilibrium Fe magnetic moments for 420 configurations are shown in Fig. 4. In the figure we observe better correlation between the DFT equilibrium magnetic moments and the ones obtained with the ensembles of mMTPs fitted to magnetic forces than with the ones without such fitting. In Tab. 3 we provide RMSEs between DFT and mMTPs equilibrium magnetic moments. As it is seen in the table, the RMSEs decrease with the increase of the number of $N_{\psi}$ for the mMTPs fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ opposed to the mMTPs fitted with $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. We note that the results are presented only for the configurations in which the magnetic moments were successfully equilibrated, i.e. when the maximal absolute magnetic force smaller than $5\cdot 10^{-3}$ meV/$\mu_{B}$ after equilibration. This quantity is referred to as equilibration reliability and is represented in Tab. 3. Indeed, all the mMTPs fitted to magnetic forces result in 100 % equilibration reliability whereas in the case when $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$ the equilibration reliability is below 100 %. A possible reason of such a result is insufficient training set size for the potentials with $N_{\psi}=3$ and $N_{\psi}=4$ that were not fitted to magnetic forces. Due to the same reason RMSE between DFT and mMTPs equilibrium magnetic moments increase with the increase of the magnetic basis.
Figure 4: Correlation between equilibrium magnetic moments of Fe calculated with mMTPs and DFT. The values of equilibrium magnetic moments obtained with the ensembles of mMTPs with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ show a better correlation with DFT than when $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$.
Table 3: RMSEs between DFT and magnetic mMTPs, equilibrium magnetic moments, and equilibration reliability obtained with different ensembles of mMTPs. We provide the values with 95% confidence interval (i.e. 2-$\sigma$ interval).
@@
To test the predictive ability of the ensembles of mMTPs depending on the number magnetic basis functions, we calculate the formation energy, lattice parameter, and total magnetic moment of the unit cell. We start from relaxation of 21 initial configurations using the mMTPs ensembles with different number of $N_{\psi}$ fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ and $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. The errors obtained during this procedure are demonstrated in Tab. 4. We also provide the relaxation reliability, which is the percent of successfully relaxed configurations (i.e. with almost zero forces, stresses and magnetic forces). As it is seen in the table, all the 21 configurations were successfully relaxed in case when mMTPs were trained with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. On the contrary, the relaxation reliability falls below 100 % when mMTPs are not fitted to magnetic forces. Thus, the training set including 2632 is sufficient when training to additional data such as magnetic forces. We also observe that the increase of the magnetic basis size significantly improves the predictive ability of the mMTPs. Finally, our results in Tab. 4 demonstrate that the mMTPs fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ yield systematically better accuracy than the mMTPs fitted with $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$ for $N_{\psi}=3$ and $N_{\psi}=4$. Due to the above reasons we provide further results only for the ensembles of mMTPs fitted to magnetic forces.
Table 4: RMSEs between DFT and mMTPs for formation energy, equilibrium lattice parameter, total cell magnetic moment per Fe atom, and the percent of successfully relaxed configurations (relaxation reliability) of 21 initial configurations. Magnetic MTPs were fitted with different magnetic basis sizes and the magnetic force weights of 0.0 $\mu_{B}^{2}$ and 0.1 $\mu_{B}^{2}$. All the RMSEs decrease with the increase of magnetic basis size. Potentials with non-zero magnetic force weight show better accuracy compared to $w_{t}=0$ for $N_{\psi}=3$ and $N_{\psi}=4$. The results for mMTPs fitted with 0.0 $\mu_{B}^{2}$ are presented only for the successfully relaxed configurations. The values are provided with 95% confidence interval (i.e. 2-$\sigma$).
@@ 2.6 Molecular dynamics with magnetic Moment Tensor Potential
For testing the mMTP at finite temperature, we conduct molecular dynamics simulations in the NPT-ensemble using the LAMMPS package [27, 28]. Lattice expansion modeling was performed for a bcc supercell of 128 atoms at 300 K. We used the most accurate ensemble of mMTPs with $N_{\psi}=4$, $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ trained on the entire dataset. The obtained trend between lattice parameter and Al concentration at 300 K is depicted and compared to the experimental data [29] in Fig. 7. Additionally, we provide the lattice parameter calculated with mMTPs and DFT at zero temperature. The supercell extension trend at finite temperature is clearly observed in comparison with the zero temperature calculations. Thus, the finite temperature lattice parameter approach the experimentally measured value. However, the RMSE between the mMTPs and the experimental lattice parameters at 300 K obtained through the “cast and quenched” approach is about 0.025 Å which is larger than the RMSE between mMTPs and DFT lattice parameters of 0.005 Å obtained at zero temperature. A possible reason is the intrinsic underestimation of the lattice parameter of bcc Fe with spin-polarized PBE calculations [30] that consequently affects the underestimation of bcc Fe-Al alloys [31]. Selection of more accurate parameters for DFT calculations or the use of other exchange-correlation potential can improve the predictive ability of the model. Nevertheless, the experimental trend is correctly captured by mMTP.
@@ Conclusion
In this paper we developed a novel method for fitting magnetic Moment Tensor Potentials (mMTPs) [16, 18]. The methodology includes fitting to magnetic forces (negative derivatives of energies with respect to magnetic moments) in addition to energies, forces, and stresses. To test the methodology we constructed a training set containing 16-atom bcc Fe-Al supercells with different concentrations of Al and Fe using density functional theory (DFT) and cDFT calculations [21] implemented in ABINIT [24]. The resulting training set included 2632 configurations. Further, we compared two mMTPs fitting methods. In the first method used in [16, 18] we fitted mMTPs only to DFT energies, forces, and stresses. The second method included fitting to magnetic forces, obtained from DFT. First, we determined an optimal weight $w_{t}$ for fitting to magnetic forces. Next, we trained the ensembles of mMTPs with different number of magnetic basis functions using the both methods. While the novel approach did not result in an observable reduction of RMSEs for energies, forces and stresses, it significantly decreased (ten times) the RMSE for magnetic forces in comparison when mMTPs is not fitted to them. Next, we observed a better agreement between DFT-calculated equilibrium magnetic moments and that obtained with the ensemble of mMTPs fitted to magnetic forces. On the contrary, mMTPs without fitting to magnetic forces showed a higher deviation from DFT-calculated values. Additionally, the mMTPs fitted to magnetic forces gave systematically better results in predicting formation energies, lattice parameters, and unit cell magnetic moments. Moreover, we verified the impact of magnetic basis size on the accuracy of prediction of bcc Fe-Al properties. Thus, we demonstrated a higher prediction accuracy when the magnetic basis size is increased. Next, we assessed the equilibration and relaxation reliability which is the percent of configurations with successfully equilibrated magnetic moments and relaxed geometry. We demonstrated that mMTPs fitted to magnetic forces always result in 100 % reliability, while mMTPs without such fitting can not successfully equilibrate and relax all the configurations from the training set. This fact indicates a lack of the data in the training set in the case when we do not fit mMTPs to magnetic forces as opposed to the novel algorithm: all of the mMTPs were able to equilibrate and relax all the configurations. Finally, we conducted molecular dynamics simulations at $T=300$ K with the ensemble of mMTPs fitted using the novel method. We correctly captured the lattice expansion trend, observed experimentally.
@@
The obtained results reveal that the presented methodology for mMTP fitting is a solution for obtaining an accurate mMTP keeping the training set size relatively small. The method may also be very effective in combination with active learning [32, 23] as it automatically selects the representative configurations for the training set, thus minimizing the number of computationally expensive spin-polarized DFT calculations.
@@
Alexey S. Kotykhov: Scripting, Calculations, Validation, Data curation, Visualization, Formal analysis, Writing - Original Draft. Konstantin Gubaev: Methodology, Software. Vadim Sotskov: Software, Formal analysis, Writing - Editing. Christian Tantardini: Scripting, Formal analysis, Writing - Editing. Max Hodapp: Resources, Calculations, Writing - Editing. Alexander V. Shapeev: Methodology, Resources, Formal analysis. Ivan S. Novikov: Methodology, Software, Formal analysis, Writing - Review & Editing.
@@ Acknowledgments
This work was supported by Russian Science Foundation (grant number 22-73-10206, https://rscf.ru/project/22-73-10206/). Moreover, M.H. gratefully acknowledges the financial support under the scope of the COMET program within the K2 Center "Integrated Computational Material, Process and Product Engineering (IC-MPPE)" (Project No 886385). This program is supported by the Austrian Federal Ministries for Climate Action, Environment, Energy, Mobility, Innovation and Technology (BMK) and for Labour and Economy (BMAW), represented by the Austrian Research Promotion Agency (FFG), and the federal states of Styria, Upper Austria and Tyrol. Ch.T. would like to thank Prof. Stefano Baroni and Prof. Stefano de Gironcoli at SISSA Trieste for useful discussion.

### v3

@@ 2.4 Correlation between DFT and magnetic Moment Tensor Potential equilibrium magnetic moments
For the additional verification of the fitted ensembles of potentials we estimate the quality of atomic magnetic moment equilibration. For this purpose we choose the configurations with the equilibrium magnetic moments from the training set, equilibrate these magnetic moments with the ensembles of mMTPs, and compare the result with DFT. The correlation between the DFT and mMTPs equilibrium Fe magnetic moments for 420 configurations are shown in Fig. 4. In the figure we observe a better correlation between the DFT equilibrium magnetic moments and the ones obtained with the ensembles of mMTPs fitted to magnetic forces than with the ones without such fitting. In Table 3 we provide RMSEs between DFT and mMTPs equilibrium magnetic moments. As seen in the table, the RMSEs decrease with the increase of the number of $N_{\psi}$ for the mMTPs fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ as opposed to the mMTPs fitted with $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. We note that the results are presented only for the configurations in which the magnetic moments were successfully equilibrated, i.e. when the maximal absolute magnetic force smaller than $5\cdot 10^{-3}$ meV/$\mu_{B}$ after equilibration. We refer to the percentage of configurations with successfully equilibrated magnetic moments as equilibration reliability and present it in Table 3. Indeed, all the mMTPs fitted to magnetic forces result in 100 % equilibration reliability whereas in the case when $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$ the equilibration reliability is below 100 %. A possible reason for such a result is insufficient training set size for the potentials with $N_{\psi}=3$ and $N_{\psi}=4$ that were not fitted to magnetic forces. Due to the same reason RMSE between DFT and mMTPs equilibrium magnetic moments increase with the increase of the magnetic basis.
Figure 4: Correlation between equilibrium magnetic moments of Fe calculated with mMTPs and DFT. Pearson correlation coefficient $r$ between mMTP and DFT magnetic moments is greater for the ensembles of mMTPs with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ than when $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$.
Table 3: RMSEs between DFT and magnetic mMTPs, equilibrium magnetic moments, and equilibration reliability obtained with different ensembles of mMTPs. We provide the values with 95% confidence interval (i.e., 2-$\sigma$ interval).
@@
To test the predictive ability of the ensembles of mMTPs depending on the number of magnetic basis functions, we calculate the formation energy, lattice parameter, and total magnetic moment of the unit cell. We start from relaxation of 21 initial configurations using the mMTPs ensembles with different number of $N_{\psi}$ fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ and $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. The errors obtained during this procedure are demonstrated in Table 4. We also provide the relaxation reliability, which is the percentage of successfully relaxed configurations (i.e., with almost zero forces, stresses and magnetic forces). As it is seen in the table, all the 21 configurations were successfully relaxed in case when mMTPs were trained with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. On the contrary, the relaxation reliability falls below 100 % when mMTPs are not fitted to magnetic forces. Thus, the training set including 2632 is sufficient when training to additional data such as magnetic forces. We also observe that the increase of the magnetic basis size significantly improves the predictive ability of the mMTPs. Finally, our results in Table 4 demonstrate that the mMTPs fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ yield a systematically better accuracy than the mMTPs fitted with $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$ for $N_{\psi}=3$ and $N_{\psi}=4$. Due to the above reasons we provide further results only for the ensembles of mMTPs fitted to magnetic forces.
Table 4: RMSEs between DFT and mMTPs for formation energy, equilibrium lattice parameter, total cell magnetic moment per Fe atom, and the percentage of successfully relaxed configurations (relaxation reliability) of 21 initial configurations. Magnetic MTPs were fitted with different magnetic basis sizes and the magnetic force weights of 0.0 $\mu_{B}^{2}$ and 0.1 $\mu_{B}^{2}$. All the RMSEs decrease with the increase of magnetic basis size. Potentials with non-zero magnetic force weight show better accuracy compared to $w_{t}=0$ for $N_{\psi}=3$ and $N_{\psi}=4$. The results for mMTPs fitted with 0.0 $\mu_{B}^{2}$ are presented only for the successfully relaxed configurations. The values are provided with 95% confidence interval (i.e., 2-$\sigma$).
@@ 2.6 Molecular dynamics with magnetic Moment Tensor Potential
In the previous sections we tested the ensembles of mMTPs for the 16-atom configurations at zero temperature. Next, we test the mMTP at finite temperature by conducting molecular dynamics (MD) simulations in the NPT-ensemble (the isotermal-isobaric ensemble in which the number of particles $N$, pressure $P$, and temperature $T$ are constant) using the LAMMPS package [31, 32]. Lattice expansion modeling was performed for a bcc supercell of 128 atoms at 300 K. This way we test both the applicability of mMTPs for studying the Fe-Al system at finite temperatures and the transferability of the potentials to configurations with a larger number of atoms none of which were not included in the training set. We used the most accurate ensemble of mMTPs with $N_{\psi}=4$, $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ trained on the entire dataset. The obtained trend between the lattice parameter and Al concentration at 300 K is plotted and compared to the experimental data [33] in Fig. 7. Additionally, we provide the lattice parameter calculated with mMTPs and DFT at zero temperature. We note that one of the five mMTPs from the ensemble demonstrated non-physical behavior, i.e., non-periodic fluctuations of the lattice parameter during MD simulations for the configurations with 43.75 % and 50 % of Al. Therefore, we excluded this potential in the calculations with these concentrations. Additionally, there was another mMTP from this ensemble that significantly underestimated the lattice parameters of configurations with 37.5 % and 43.75 % concentration of Al. We did not exclude this potential which resulted in clearly observed increase of the confidence intervals in Fig. 7 for these concentrations of Al. The reason behind non-physical behavior of one of the mMTPs and the underestimation of the lattice parameters with the other potential is the problem with transferability from the 16-atom systems, used in the training set, to the system including 128 atoms, that were not used during the fitting. Therefore, to model such systems one can rely either on manually testing and selecting suitable potentials, as we have done in this work, or on the use of the active learning algorithm. This algorithm estimates the extrapolation grade of configurations “on-the-fly” and selects the most representative configurations for mMTP re-fitting (see, e.g. [34]). The supercell extension trend at finite temperature is clearly observed in Fig. 7 in comparison with the zero temperature calculations. Thus, the finite-temperature lattice parameter approaches the experimentally measured value. However, the RMSE between the mMTPs and the experimental lattice parameters at 300 K obtained through the “cast and quenched” approach is about 0.025 Å which is larger than the RMSE between mMTPs and DFT lattice parameters of 0.005 Å obtained at zero temperature. A possible reason for the underestimation of the lattice parameter of bcc Fe is the Perdew-Burke-Ernzerhof DFT functional [35], that is reported to affect the underestimation of the lattice parameter for bcc Fe-Al alloys [36]. Consequently, this may also lead to errors in prediction of magnetic properties, e.g., total magnetic moment of the unit cell as it depends on the cell volume (see, e.g., [18]). Selection of more accurate parameters for DFT calculations or the use of other exchange-correlation potential can improve the predictive ability of both spin-polarized DFT calculation and the ensemble of mMTPs fitted to the results of spin-polarized DFT. Nevertheless, the experimental trend is correctly captured by mMTP.
@@
2.7 Accelerating DFT calculations with magnetic Moment Tensor Potential
To compare the computational cost of spin-polarized DFT and mMTP we estimate the average computational time needed to calculate energy, forces, stresses, and equilibrium magnetic moments of several hundred different 16-atom Fe-Al cofigurations. Average calculation time on a single central processing unit (CPU) core is about $10^{4}$ s/atom for spin-polarized DFT and approximately $10^{-2}$ s/atom for mMTP with $N_{\psi}=4$. Therefore, mMTP is several orders of magnitude faster than spin-polarized DFT. This fact clearly demonstrates the relevance of developing magnetic MLIPs for such kind of simulations.
@@ Conclusion
In this paper we developed a method for fitting magnetic Moment Tensor Potentials (mMTPs) [18, 20] to magnetic forces (negative derivatives of energies with respect to magnetic moments) in addition to energies, forces, and stresses. To test the methodology we constructed a training set containing 16-atom bcc Fe-Al supercells with different concentrations of Al and Fe using density functional theory (DFT) and cDFT calculations [24] implemented in ABINIT [28]. The resulting training set included 2632 configurations. Further, we compared two mMTPs fitting methods. In the first method used in [18, 20] we fitted mMTPs only to DFT energies, forces, and stresses. The second method included fitting to magnetic forces, obtained from DFT. First, we determined an optimal weight $w_{t}$ for fitting to magnetic forces. Next, we trained the ensembles of mMTPs with different number of magnetic basis functions using the both methods. While the our developed approach did not result in an observable reduction of RMSEs for energies, forces and stresses, it significantly decreased (ten times) the RMSE for magnetic forces in comparison when mMTPs is not fitted to them. Next, we observed a better agreement between DFT-calculated equilibrium magnetic moments and those obtained with the ensemble of mMTPs fitted to magnetic forces. On the contrary, mMTPs without fitting to magnetic forces showed a higher deviation from DFT-calculated values. Additionally, the mMTPs fitted to magnetic forces gave systematically better results in predicting formation energies, lattice parameters, and unit cell magnetic moments. Moreover, we verified the impact of magnetic basis size on the accuracy of prediction of bcc Fe-Al properties. Thus, we demonstrated a higher prediction accuracy when the magnetic basis size is increased. Next, we assessed the equilibration and relaxation reliability which is the percentage of configurations with successfully equilibrated magnetic moments and relaxed geometry. We demonstrated that mMTPs fitted to magnetic forces always result in 100 % reliability, while mMTPs without such fitting can not successfully equilibrate and relax all the configurations from the training set. This fact indicates a lack of the data in the training set in the case when we do not fit mMTPs to magnetic forces as opposed to the novel algorithm: all of the mMTPs were able to equilibrate and relax all the configurations. Finally, we conducted molecular dynamics simulations at $T=300$ K with the ensemble of mMTPs fitted using the novel method. We correctly captured the experimentally observed lattice expansion trend.
@@
The obtained results reveal that the presented methodology for mMTP fitting is a tool for obtaining an accurate mMTP keeping the training set size relatively small. The method should also be very effective in combination with active learning [34, 26] as it automatically selects the representative configurations for the training set, thus minimizing the number of computationally expensive spin-polarized DFT calculations.
@@
Alexey S. Kotykhov: Scripting, Calculations, Validation, Data curation, Visualization, Formal analysis, Writing - Original Draft. Konstantin Gubaev: Methodology, Software. Vadim Sotskov: Software, Formal analysis, Writing - Editing. Christian Tantardini: Scripting, Formal analysis, Writing - Editing. Max Hodapp: Resources, Calculations, Writing - Editing. Alexander V. Shapeev: Methodology, Resources, Formal analysis, Writing - Review & Editing. Ivan S. Novikov: Methodology, Software, Formal analysis, Writing - Review & Editing.
@@ Acknowledgments
This work was supported by Russian Science Foundation (grant number 22-73-10206, https://rscf.ru/project/22-73-10206/). Moreover, M.H. gratefully acknowledges the financial support under the scope of the COMET program within the K2 Center "Integrated Computational Material, Process and Product Engineering (IC-MPPE)" (Project No 886385). This program is supported by the Austrian Federal Ministries for Climate Action, Environment, Energy, Mobility, Innovation and Technology (BMK) and for Labour and Economy (BMAW), represented by the Austrian Research Promotion Agency (FFG), and the federal states of Styria, Upper Austria and Tyrol. Ch.T. would like to thank Prof. Stefano Baroni and Prof. Stefano de Gironcoli at SISSA Trieste for useful discussion. Ch.T was supported by the Norwegian Research Council through a Centre of Excellence grant (Hylleraas Centre 262695), a FRIPRO grant (ReMRChem 324590).
Data availability
ll the ensembles of the fitted potentials and the training set are available in the https://gitlab.com/ivannovikov/datasets_for_magnetic_MTP repository in the Fe_Al_fitting_to_magnetic_forces/ folder.